### v4

@@ 5.2 Numerical Results
We aim to compare numerically the asymptotic error distributions of the schemes introduced in Sections 14. We fix a small prescribed accuracy $\varepsilon\in(0,1)$, set $h$ as in Theorems 1.8 and 2.8 for the NSA and ANSA schemes and set $L$ and $\mathbf{N}=\{N_{\ell},0\leq\ell\leq L\}$ as in Theorems 3.8 and 4.4 for the MLSA and AMLSA schemes. We run each scheme $5000$ times and plot the corresponding joint distributions of the renormalized $(\text{VaR},\text{ES})$ estimation errors as in Corollaries 1.4 and 2.4 and Theorems 3.5 and 4.1. The unbiased SA and averaged SA (ASA) schemes of [8, Theorem 3.4] and [2, Theorem 2.4], that are based on simulating the loss $X_{0}$ exactly via (32), are also run for benchmarking purposes. But, unlike [2, Theorem 2.4], we do not average out the ES component. Qualitatively however, the resulting outputs should be comparable.
For the case study, we set $S_{0}=1$, $r=2\%$, $\kappa=12\%$, $\sigma=20\%$, $T=1\,\mathrm{year}$, $\Delta T_{i}=3\,\mathrm{months}$, $\delta=7\;\mathrm{days}$ and $\alpha=85\%$. We use a $30/360$ day count fraction convention. (33) and (34) yield $\xi^{0}_{\star}\approx 2.19$ and $\chi^{0}_{\star}\approx 3.29$. The biased risk measures $\xi^{h_{L}}_{\star}$ and $\chi^{h_{L}}_{\star}$ needed for the CLTs of Theorems 3.5 and 4.1 are computed by averaging out $200$ outcomes of the NSA scheme with bias $h=h_{L}$ and $n=10^{5}$ iterations, yielding $\xi^{h_{L}}_{\star}\approx 2.17$ and $\chi^{h_{L}}_{\star}\approx 3.41$. We set $\varepsilon=\frac{1}{256}$ and $\beta=0.9$ for all six SA schemes. The choice of $\beta$ is such that it is in the interval $\big(\frac{8}{9},1\big)$ for which Theorem 4.1 is proven. It is equal for all schemes to allow comparison. We adopt $\{\gamma_{n}=1\times n^{-0.9},n\geq 1\}$ for the unbiased SA and ASA algorithms, $\{\gamma_{n}=0.1\times(250+n)^{-0.9},n\geq 1\}$ for the NSA and ANSA algorithms, and $h_{0}=\frac{1}{32}$, $M=2$, $L=3$ and $\{\gamma_{n}=0.1\times(1500+n)^{-0.9},n\geq 1\}$ for the MLSA and AMLSA algorithms. We also compute the theoretical ES variance factors given in the CLTs via the Monte Carlo formulas in Remarks 1.6(iv) and 3.7(vii).
@@ Conclusion
All things considered, the analyses in this paper hint at resorting to a multilevel scheme rather than a simply nested one in order to reduce complexity, and at applying an averaged scheme rather than a non-averaged one in order to increase numerical stability. For some prescribed accuracy $\varepsilon>0$, the optimal complexity attained by the presented multilevel algorithms is in $\mathrm{O}(\varepsilon^{-\frac{5}{2}})$. Given a learning rate $\{\gamma_{n}=\gamma_{1}n^{-\beta},n\geq 1\}$, $\gamma_{1}>0$, $\beta\in\big(\frac{1}{2},1\big]$, this complexity is achieved by MLSA for $\beta=1$ under the constraint $\lambda\gamma_{1}>1$, where the constant $\lambda>0$ in (9) is explicit but tedious to compute. This complexity is also achieved by AMLSA for $\beta\in\big(\frac{8}{9},1\big)$, without any constraint on $\gamma_{1}$.
However, $\mathrm{O}(\varepsilon^{-\frac{5}{2}})$ remains higher than the theoretical optimum of $\mathrm{O}(\varepsilon^{-2})$ that multilevel algorithms are known for [18, 21]. This gap in performance stems from the discontinuity of the gradient used in the updating formula (7), registering an $\mathrm{O}(1)$ error whenever the generated loss $X_{h}$ is too close to the estimate $\xi^{h}_{n}$ but falls on the opposite side of the discontinuity with respect to the simulation target $X_{0}$ [26, 23]. This limitation should be addressed in future research.
@@ Appendix A Auxiliary Results
    $\limsup_{n\rightarrow\infty}\gamma_{n}^{-b}\sum_{k=1}^{n}\gamma_{k}^{1+b}e_{k}\exp\bigg(-\lambda\sum_{j=k+1}^{n}\gamma_{j}\bigg)\leq\frac{1}{C}\limsup_{n\rightarrow\infty}e_{n},$
@@
    $\limsup_{n\rightarrow\infty}{\gamma_{n}^{-b}\exp\bigg(-\lambda\sum_{j=1}^{n}\gamma_{j}\bigg)}=0.$
@@
    $\xi^{h}_{\star}-\xi^{0}_{\star}=-\frac{v(\xi^{0}_{\star})}{f_{X_{0}}(\xi^{0}_{\star})}h+\mathrm{o}(h),\qquad\chi^{h}_{\star}-\chi^{0}_{\star}=-h\int_{\xi^{0}_{\star}}^{\infty}\frac{v(\xi)}{1-\alpha}\mathrm{d}\xi+\mathrm{o}(h).$
@@
    Suppose that $\varphi(Y,Z)\in L^{2}(\mathbb{P})$ and that Assumptions 1.1 and 1.1 hold. If $\gamma_{n}=\gamma_{1}n^{-\beta}$, $n\geq 1$, $\gamma_{1}>0$, $\beta\in(0,1]$, with $\lambda\gamma_{1}>1$ if $\beta=1$, then there exists a positive constant $C<\infty$ such that, for any positive integer $n$,
@@
The sequence $\{\xi^{h}_{n},n\geq 0\}$ of dynamics (7) decomposes into
$\xi^{h}_{n}-\xi^{h}_{\star}=\big(1-\gamma_{n}V_{0}^{\prime\prime}(\xi^{0}_{\star})\big)(\xi^{h}_{n-1}-\xi^{h}_{\star})+\gamma_{n}g^{h}_{n}+\gamma_{n}r^{h}_{n}+\gamma_{n}\rho^{h}_{n}+\gamma_{n}e^{h}_{n},$ (35)
@@
$g^{h}_{n}$ $=\big(V_{0}^{\prime\prime}(\xi^{0}_{\star})-V_{h}^{\prime\prime}(\xi^{h}_{\star})\big)(\xi^{h}_{n-1}-\xi^{h}_{\star}),$ (36)
$r^{h}_{n}$ $=V_{h}^{\prime\prime}(\xi^{h}_{\star})(\xi^{h}_{n-1}-\xi^{h}_{\star})-V_{h}^{\prime}(\xi^{h}_{n-1}),$ (37)
$\rho^{h}_{n}$ $=V_{h}^{\prime}(\xi^{h}_{n-1})-V_{h}^{\prime}(\xi^{h}_{\star})-\big(H_{1}(\xi^{h}_{n-1},X_{h}^{(n)})-H_{1}(\xi^{h}_{\star},X_{h}^{(n)})\big),$ (38)
$e^{h}_{n}$ $=V_{h}^{\prime}(\xi^{h}_{\star})-H_{1}(\xi^{h}_{\star},X_{h}^{(n)})=-H_{1}(\xi^{h}_{\star},X_{h}^{(n)}).$ (39)
Hence, unrolling (35),
$\xi^{h}_{n}-\xi^{h}_{\star}=(\xi^{h}_{0}$ $-\xi^{h}_{\star})\Pi_{1:n}+\sum_{k=1}^{n}\gamma_{k}\Pi_{k+1:n}g^{h}_{k}$ (40)
@@
$\Pi_{k:n}=\prod_{j=k}^{n}\big(1-\gamma_{j}V_{0}^{\prime\prime}(\xi^{0}_{\star})\big),$ (41)
@@
Likewise, the sequence $\{\chi^{h}_{n},n\geq 0\}$ of dynamics (8) rewrites
$\chi^{h}_{n}-\chi^{h}_{\star}$ $=\frac{1}{n}\sum_{k=1}^{n}\Big(\xi^{h}_{k-1}+\frac{1}{1-\alpha}(X_{h}^{(k)}-\xi^{h}_{k-1})^{+}\Big)-V_{h}(\xi^{h}_{\star})$ (42)
@@
$\theta^{h}_{k}$ $=\xi_{k-1}^{h}-\xi^{h}_{\star}+\frac{1}{1-\alpha}\big((X_{h}^{(k)}-\xi_{k-1}^{h})^{+}-(X_{h}^{(k)}-\xi^{h}_{\star})^{+}\big)-\big(V_{h}(\xi_{k-1}^{h})-V_{h}(\xi^{h}_{\star})\big),$ (43)
@@
$\eta^{h}_{k}$ $=\frac{1}{1-\alpha}\big((X_{h}^{(k)}-\xi^{h}_{\star})^{+}-\mathbb{E}[(X_{h}^{(k)}-\xi^{h}_{\star})^{+}]\big).$ (44)
We first provide a useful upper bound on $\Pi_{k:n}$. Since $\lim_{k\rightarrow\infty}{\gamma_{k}}=0$, there exists $k_{0}\geq 0$ such that $(1-\gamma_{j}V_{0}^{\prime\prime}(\xi^{0}_{\star}))>0$, $j\geq k_{0}$. Thus, using the inequality $1+x\leq\exp(x)$, $x\in\mathbb{R}$, we obtain that for $n$ large enough,
$|\Pi_{k:n}|$ $=|\Pi_{k:k_{0}-1}|\prod_{j=k_{0}\vee k}^{n}\big(1-\gamma_{j}V_{0}^{\prime\prime}(\xi^{0}_{\star})\big)$ (45)
@@
Set $n=\lceil h^{-2}\rceil$. We study below the contribution of each term on the right hand sides of (40) and (42) to the asymptotic estimation error.
@@
By assumption and via (9),
$2\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})\geq 2\gamma_{1}\inf_{h\in\mathcal{H}\cup\{0\}}V_{h}^{\prime\prime}(\xi^{h}_{\star})\geq\frac{16}{3}\gamma_{1}\inf_{h\in\mathcal{H}\cup\{0\}}\lambda_{h}\geq\gamma_{1}\lambda>1\quad\mbox{ if }\quad\beta=1.$ (46)
We deduce via the inequality $\gamma_{1}^{\frac{1}{2}}\gamma_{\lceil h^{-2}\rceil}^{-\frac{1}{2}}\geq h^{-\beta}$, (46), (45) and Lemma A.1(ii) that
@@
$h^{-\beta}(\xi^{h}_{0}-\xi^{h}_{\star})\Pi_{1:\lceil h^{-2}\rceil}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
Recalling that $h^{-\beta}\leq\gamma_{1}^{\frac{1}{2}}\gamma_{\lceil h^{-2}\rceil}^{-\frac{1}{2}}$ and that $2\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})>1$ if $\beta=1$ by (46), Lemma A.1(i) yields
@@
$h^{-\beta}\sum_{k=1}^{\lceil h^{-2}\rceil}\gamma_{k}\Pi_{k+1:\lceil h^{-2}\rceil}g^{h}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$V_{h}^{\prime}(\xi)=V_{h}^{\prime\prime}(\xi^{h}_{\star})(\xi-\xi^{h}_{\star})+(\xi-\xi^{h}_{\star})\int_{0}^{1}\big(V_{h}^{\prime\prime}\big(\xi^{h}_{\star}+t(\xi-\xi^{h}_{\star})\big)-V_{h}^{\prime\prime}(\xi^{h}_{\star})\big)\mathrm{d}t.$ (47)
The uniform Lipschitz regularity of $\{V_{h}^{\prime\prime}=(1-\alpha)^{-1}f_{X_{h}},h\in\mathcal{H}\}$, by Assumption 1.1(iv), and (47) yield
@@
$\mathbb{E}[|r^{h}_{k}|]\leq C\gamma_{k}.$ (48)
Recalling that $2\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})>1$ if $\beta=1$ by (46), using (45) and Lemma A.1(i),
@@
$h^{-\beta}\sum_{k=1}^{\lceil h^{-2}\rceil}\gamma_{k}\Pi_{k+1:\lceil h^{-2}\rceil}r^{h}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$\mathbb{E}[|\rho^{h}_{k}|^{2}]$ $=\mathbb{E}\big[\mathbb{V}\mathrm{ar}\big(H_{1}(\xi^{h}_{k-1},X_{h}^{(k)})-H_{1}(\xi^{h}_{\star},X_{h}^{(k)})\big|\mathcal{F}^{h}_{k-1}\big)\big]$
@@
$\mathbb{E}[|\rho^{h}_{k}|^{2}]\leq\mathbb{E}[(H_{1}(\xi^{h}_{k-1},X_{h}^{(k)})-H_{1}(\xi^{h}_{\star},X_{h}^{(k)}))^{2}]\leq C\gamma_{k}^{\frac{1}{2}}.$ (49)
By (6) and (38), $\mathbb{E}[\rho^{h}_{k}|\mathcal{F}_{k-1}^{h}]=0$, so that $\{\rho^{h}_{k},k\geq 1\}$ are $\{\mathcal{F}^{h}_{k},k\geq 1\}$-martingale increments. Hence, via (49),
@@
Since $h^{-\beta}\leq\gamma_{1}^{\frac{1}{2}}\gamma_{\lceil h^{-2}\rceil}^{-\frac{1}{2}}$ and $2\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})>1$ if $\beta=1$ by (46), via (45) and Lemma A.1(i),
@@
$h^{-\beta}\sum_{k=1}^{\lceil h^{-2}\rceil}\gamma_{k}\Pi_{k+1:\lceil h^{-2}\rceil}\rho^{h}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
Via (4) and (43), $\mathbb{E}[\theta_{k}^{h}|\mathcal{F}_{k-1}^{h}]=0$, i.e. $\{\theta^{h}_{k},k\geq 1\}$ are $\{\mathcal{F}^{h}_{k},k\geq 1\}$-martingale increments. Hence, by Lemma A.3(ii),
$\mathbb{E}\bigg[\bigg(\frac{1}{\sqrt{n}}\sum_{k=1}^{n}\theta^{h}_{k}\bigg)^{2}\bigg]$ $=\frac{1}{n}\sum_{k=1}^{n}\mathbb{E}[|\theta^{h}_{k}|^{2}]$ (50)
@@
$\frac{h^{-1}}{\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}\theta^{h}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$\mathbb{E}\bigg[\bigg|\frac{1}{\sqrt{n}}\sum_{k=1}^{n}\zeta^{h}_{k}\bigg|\bigg]$ $\leq\frac{\sup_{h^{\prime}\in\mathcal{H}}\|f_{X_{h^{\prime}}}\|_{\infty}}{2(1-\alpha)}\bigg(\frac{1}{\sqrt{n}}\,\mathbb{E}[(\xi^{h}_{0}-\xi^{h}_{\star})^{2}]+\frac{1}{\sqrt{n}}\sum_{k=2}^{n}\mathbb{E}[(\xi_{k-1}^{h}-\xi^{h}_{\star})^{2}]\bigg)$ (51)
@@
$\frac{h^{-1}}{\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}\zeta^{h}_{k}\mathrel{{\mathop{\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$

### v5

@@ 5.2 Numerical Results
We aim to compare numerically the asymptotic error distributions of the schemes introduced in Sections 14. We fix a small prescribed accuracy $\varepsilon\in(0,1)$, set $h$ as in Theorems 1.8 and 2.6 for the NSA and ANSA schemes and set $L$ and $\mathbf{N}=\{N_{\ell},0\leq\ell\leq L\}$ as in Theorems 3.7 and 4.3 for the MLSA and AMLSA schemes. We run each scheme $5000$ times and plot the corresponding joint distributions of the renormalized $(\text{VaR},\text{ES})$ estimation errors as in Corollaries 1.5 and 2.3 and Theorems 3.5 and 4.1. The unbiased SA and averaged SA (ASA) schemes of [8, Theorem 3.4] and [2, Theorem 2.4], that are based on simulating the loss $X_{0}$ exactly via (5.3), are also run for benchmarking purposes. But, unlike [2, Theorem 2.4], we do not average out the ES component. Qualitatively however, the resulting outputs should be comparable.
For the case study, we set $S_{0}=1$, $r=2\%$, $\kappa=12\%$, $\sigma=20\%$, $T=1\,\mathrm{year}$, $\Delta T_{i}=3\,\mathrm{months}$, $\delta=7\;\mathrm{days}$ and $\alpha=85\%$. We use a $30/360$ day count fraction convention. (5.4) and (5.5) yield $\xi^{0}_{\star}\approx 2.19$ and $\chi^{0}_{\star}\approx 3.29$. The biased risk measures $\xi^{h_{L}}_{\star}$ and $\chi^{h_{L}}_{\star}$ needed for the CLTs of Theorems 3.5 and 4.1 are computed by averaging out $200$ outcomes of the NSA scheme with bias $h=h_{L}$ and $n=10^{5}$ iterations, yielding $\xi^{h_{L}}_{\star}\approx 2.17$ and $\chi^{h_{L}}_{\star}\approx 3.41$. We set $\varepsilon=\frac{1}{256}$ and $\beta=0.9$ for all six SA schemes. The choice of $\beta$ is such that it is in the interval $\big(\frac{8}{9},1\big)$ for which Theorem 4.1 is proven. It is equal for all schemes to allow comparison. We adopt $\{\gamma_{n}=1\times n^{-0.9},n\geq 1\}$ for the unbiased SA and ASA algorithms, $\{\gamma_{n}=0.1\times(250+n)^{-0.9},n\geq 1\}$ for the NSA and ANSA algorithms, and $h_{0}=\frac{1}{32}$, $M=2$, $L=3$ and $\{\gamma_{n}=0.1\times(1500+n)^{-0.9},n\geq 1\}$ for the MLSA and AMLSA algorithms. We also compute the theoretical ES variance factors given in the CLTs via the Monte Carlo formulas in Remarks 1.6(iv) and 3.6(vii).
@@ Conclusion
All things considered, the analyses in this paper hint at resorting to a multilevel scheme rather than a simply nested one in order to reduce complexity, and at applying an averaged scheme rather than a non-averaged one in order to increase numerical stability. For some prescribed accuracy $\varepsilon>0$, the optimal complexity attained by the presented multilevel algorithms is in $\operatorname{O}(\varepsilon^{-\frac{5}{2}})$. Given a learning rate $\{\gamma_{n}=\gamma_{1}n^{-\beta},n\geq 1\}$, $\gamma_{1}>0$, $\beta\in\big(\frac{1}{2},1\big]$, this complexity is achieved by MLSA for $\beta=1$ under the constraint $\lambda\gamma_{1}>1$, where the constant $\lambda>0$ in (1.6) is explicit but tedious to compute. This complexity is also achieved by AMLSA for $\beta\in\big(\frac{8}{9},1\big)$, without any constraint on $\gamma_{1}$.
However, $\operatorname{O}(\varepsilon^{-\frac{5}{2}})$ remains higher than the theoretical optimum of $\operatorname{O}(\varepsilon^{-2})$ that multilevel algorithms are known for [18, 21]. This gap in performance stems from the discontinuity of the gradient used in the updating formula (1.4), registering an $\operatorname{O}(1)$ error whenever the generated loss $X_{h}$ is too close to the estimate $\xi^{h}_{n}$ but falls on the opposite side of the discontinuity with respect to the simulation target $X_{0}$ [26, 23]. This limitation should be addressed in future research.
@@ Appendix A Auxiliary Results
    $\limsup_{n\to\infty}\gamma_{n}^{-b}\sum_{k=1}^{n}\gamma_{k}^{1+b}e_{k}\exp\bigg(-\lambda\sum_{j=k+1}^{n}\gamma_{j}\bigg)\leq\frac{1}{C}\limsup_{n\to\infty}e_{n},$
@@
    $\limsup_{n\to\infty}{\gamma_{n}^{-b}\exp\bigg(-\lambda\sum_{j=1}^{n}\gamma_{j}\bigg)}=0.$
@@
    $\xi^{h}_{\star}-\xi^{0}_{\star}=-\frac{v(\xi^{0}_{\star})}{f_{X_{0}}(\xi^{0}_{\star})}h+\operatorname{o}(h),\qquad\chi^{h}_{\star}-\chi^{0}_{\star}=-h\int_{\xi^{0}_{\star}}^{\infty}\frac{v(\xi)}{1-\alpha}\mathrm{d}\xi+\operatorname{o}(h).$
@@
    Suppose that $\varphi(Y,Z)\in L^{2}(\mathbb{P})$ and that Assumptions 1.1 and 1.3 hold. If $\gamma_{n}=\gamma_{1}n^{-\beta}$, $n\geq 1$, $\gamma_{1}>0$, $\beta\in(0,1]$, with $\lambda\gamma_{1}>1$ if $\beta=1$, then there exists a positive constant $C<\infty$ such that, for any positive integer $n$,
@@
The sequence $\{\xi^{h}_{n},n\geq 0\}$ of dynamics (1.4) decomposes into
$\xi^{h}_{n}-\xi^{h}_{\star}=\big(1-\gamma_{n}V_{0}^{\prime\prime}(\xi^{0}_{\star})\big)(\xi^{h}_{n-1}-\xi^{h}_{\star})+\gamma_{n}g^{h}_{n}+\gamma_{n}r^{h}_{n}+\gamma_{n}\rho^{h}_{n}+\gamma_{n}e^{h}_{n},$ (B.1)
@@
$g^{h}_{n}$ $=\big(V_{0}^{\prime\prime}(\xi^{0}_{\star})-V_{h}^{\prime\prime}(\xi^{h}_{\star})\big)(\xi^{h}_{n-1}-\xi^{h}_{\star}),$ (B.2)
$r^{h}_{n}$ $=V_{h}^{\prime\prime}(\xi^{h}_{\star})(\xi^{h}_{n-1}-\xi^{h}_{\star})-V_{h}^{\prime}(\xi^{h}_{n-1}),$ (B.3)
$\rho^{h}_{n}$ $=V_{h}^{\prime}(\xi^{h}_{n-1})-V_{h}^{\prime}(\xi^{h}_{\star})-\big(H_{1}(\xi^{h}_{n-1},X_{h}^{(n)})-H_{1}(\xi^{h}_{\star},X_{h}^{(n)})\big),$ (B.4)
$e^{h}_{n}$ $=V_{h}^{\prime}(\xi^{h}_{\star})-H_{1}(\xi^{h}_{\star},X_{h}^{(n)})=-H_{1}(\xi^{h}_{\star},X_{h}^{(n)}).$ (B.5)
Hence, unrolling (B.1),
$\xi^{h}_{n}-\xi^{h}_{\star}=(\xi^{h}_{0}$ $-\xi^{h}_{\star})\Pi_{1:n}+\sum_{k=1}^{n}\gamma_{k}\Pi_{k+1:n}g^{h}_{k}$ (B.6)
@@
$\Pi_{k:n}=\prod_{j=k}^{n}\big(1-\gamma_{j}V_{0}^{\prime\prime}(\xi^{0}_{\star})\big),$ (B.7)
@@
Likewise, the sequence $\{\chi^{h}_{n},n\geq 0\}$ of dynamics (1.5) rewrites
$\chi^{h}_{n}-\chi^{h}_{\star}$ $=\frac{1}{n}\sum_{k=1}^{n}\Big(\xi^{h}_{k-1}+\frac{1}{1-\alpha}(X_{h}^{(k)}-\xi^{h}_{k-1})^{+}\Big)-V_{h}(\xi^{h}_{\star})$ (B.8)
@@
$\theta^{h}_{k}$ $=\xi_{k-1}^{h}-\xi^{h}_{\star}+\frac{1}{1-\alpha}\big((X_{h}^{(k)}-\xi_{k-1}^{h})^{+}-(X_{h}^{(k)}-\xi^{h}_{\star})^{+}\big)-\big(V_{h}(\xi_{k-1}^{h})-V_{h}(\xi^{h}_{\star})\big),$ (B.9)
@@
$\eta^{h}_{k}$ $=\frac{1}{1-\alpha}\big((X_{h}^{(k)}-\xi^{h}_{\star})^{+}-\mathbb{E}[(X_{h}^{(k)}-\xi^{h}_{\star})^{+}]\big).$ (B.10)
We first provide a useful upper bound on $\Pi_{k:n}$. Since $\lim_{k\to\infty}{\gamma_{k}}=0$, there exists $k_{0}\geq 0$ such that $(1-\gamma_{j}V_{0}^{\prime\prime}(\xi^{0}_{\star}))>0$, $j\geq k_{0}$. Thus, using the inequality $1+x\leq\exp(x)$, $x\in\mathbb{R}$, we obtain that for $n$ large enough,
$|\Pi_{k:n}|$ $=|\Pi_{k:k_{0}-1}|\prod_{j=k_{0}\vee k}^{n}\big(1-\gamma_{j}V_{0}^{\prime\prime}(\xi^{0}_{\star})\big)$ (B.11)
@@
Set $n=\lceil h^{-2}\rceil$. We study below the contribution of each term on the right hand sides of (B.6) and (B.8) to the asymptotic estimation error.
@@
By assumption and via (1.6),
$2\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})\geq 2\gamma_{1}\inf_{h\in\mathcal{H}\cup\{0\}}V_{h}^{\prime\prime}(\xi^{h}_{\star})\geq\frac{16}{3}\gamma_{1}\inf_{h\in\mathcal{H}\cup\{0\}}\lambda_{h}\geq\gamma_{1}\lambda>1\quad\mbox{ if }\quad\beta=1.$ (B.12)
We deduce via the inequality $\gamma_{1}^{\frac{1}{2}}\gamma_{\lceil h^{-2}\rceil}^{-\frac{1}{2}}\geq h^{-\beta}$, (B.12), (B.11) and Lemma A.1(ii) that
@@
$h^{-\beta}(\xi^{h}_{0}-\xi^{h}_{\star})\Pi_{1:\lceil h^{-2}\rceil}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
Recalling that $h^{-\beta}\leq\gamma_{1}^{\frac{1}{2}}\gamma_{\lceil h^{-2}\rceil}^{-\frac{1}{2}}$ and that $2\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})>1$ if $\beta=1$ by (B.12), Lemma A.1(i) yields
@@
$h^{-\beta}\sum_{k=1}^{\lceil h^{-2}\rceil}\gamma_{k}\Pi_{k+1:\lceil h^{-2}\rceil}g^{h}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$V_{h}^{\prime}(\xi)=V_{h}^{\prime\prime}(\xi^{h}_{\star})(\xi-\xi^{h}_{\star})+(\xi-\xi^{h}_{\star})\int_{0}^{1}\big(V_{h}^{\prime\prime}\big(\xi^{h}_{\star}+t(\xi-\xi^{h}_{\star})\big)-V_{h}^{\prime\prime}(\xi^{h}_{\star})\big)\mathrm{d}t.$ (B.13)
The uniform Lipschitz regularity of $\{V_{h}^{\prime\prime}=(1-\alpha)^{-1}f_{X_{h}},h\in\mathcal{H}\}$, by Assumption 1.1(iv), and (B.13) yield
@@
$\mathbb{E}[|r^{h}_{k}|]\leq C\gamma_{k}.$ (B.14)
Recalling that $2\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})>1$ if $\beta=1$ by (B.12), using (B.11) and Lemma A.1(i),
@@
$h^{-\beta}\sum_{k=1}^{\lceil h^{-2}\rceil}\gamma_{k}\Pi_{k+1:\lceil h^{-2}\rceil}r^{h}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$\mathbb{E}[|\rho^{h}_{k}|^{2}]$ $=\mathbb{E}\big[\operatorname{\mathbb{V}ar}\big(H_{1}(\xi^{h}_{k-1},X_{h}^{(k)})-H_{1}(\xi^{h}_{\star},X_{h}^{(k)})\big|\mathcal{F}^{h}_{k-1}\big)\big]$
@@
$\mathbb{E}[|\rho^{h}_{k}|^{2}]\leq\mathbb{E}[(H_{1}(\xi^{h}_{k-1},X_{h}^{(k)})-H_{1}(\xi^{h}_{\star},X_{h}^{(k)}))^{2}]\leq C\gamma_{k}^{\frac{1}{2}}.$ (B.15)
By (1.3) and (B.4), $\mathbb{E}[\rho^{h}_{k}|\mathcal{F}_{k-1}^{h}]=0$, so that $\{\rho^{h}_{k},k\geq 1\}$ are $\{\mathcal{F}^{h}_{k},k\geq 1\}$-martingale increments. Hence, via (B.15),
@@
Since $h^{-\beta}\leq\gamma_{1}^{\frac{1}{2}}\gamma_{\lceil h^{-2}\rceil}^{-\frac{1}{2}}$ and $2\gamma_{1}V_{0}^{\prime\prime}(\xi^{0}_{\star})>1$ if $\beta=1$ by (B.12), via (B.11) and Lemma A.1(i),
@@
$h^{-\beta}\sum_{k=1}^{\lceil h^{-2}\rceil}\gamma_{k}\Pi_{k+1:\lceil h^{-2}\rceil}\rho^{h}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
Via (1.1) and (B.9), $\mathbb{E}[\theta_{k}^{h}|\mathcal{F}_{k-1}^{h}]=0$, i.e. $\{\theta^{h}_{k},k\geq 1\}$ are $\{\mathcal{F}^{h}_{k},k\geq 1\}$-martingale increments. Hence, by Lemma A.3(ii),
$\mathbb{E}\bigg[\bigg(\frac{1}{\sqrt{n}}\sum_{k=1}^{n}\theta^{h}_{k}\bigg)^{2}\bigg]$ $=\frac{1}{n}\sum_{k=1}^{n}\mathbb{E}[|\theta^{h}_{k}|^{2}]$ (B.16)
@@
$\frac{h^{-1}}{\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}\theta^{h}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{2}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$
@@
$\mathbb{E}\bigg[\bigg|\frac{1}{\sqrt{n}}\sum_{k=1}^{n}\zeta^{h}_{k}\bigg|\bigg]$ $\leq\frac{\sup_{h^{\prime}\in\mathcal{H}}\|f_{X_{h^{\prime}}}\|_{\infty}}{2(1-\alpha)}\bigg(\frac{1}{\sqrt{n}}\,\mathbb{E}[(\xi^{h}_{0}-\xi^{h}_{\star})^{2}]+\frac{1}{\sqrt{n}}\sum_{k=2}^{n}\mathbb{E}[(\xi_{k-1}^{h}-\xi^{h}_{\star})^{2}]\bigg)$ (B.17)
@@
$\frac{h^{-1}}{\lceil h^{-2}\rceil}\sum_{k=1}^{\lceil h^{-2}\rceil}\zeta^{h}_{k}\mathrel{{\mathop{\longrightarrow}\limits^{L^{1}(\mathbb{P})}}}0\quad\mbox{ as }\quad\mathcal{H}\ni h\downarrow 0.$